# unPIC: A Geometric Multiview Prior for Image to 3D Synthesis

### v2

@@ Abstract
We introduce a hierarchical probabilistic approach to go from a 2D image to multiview 3D: a diffusion “prior” predicts the unseen 3D geometry, which then conditions a diffusion “decoder” to generate novel views of the subject. We use a pointmap-based geometric representation to coordinate the generation of multiple target views simultaneously. We construct a predictable distribution of geometric features per target view to enable learnability across examples, and generalization to arbitrary inputs images. Our modular, geometry-driven approach to novel-view synthesis (called “unPIC”) beats competing baselines such as CAT3D, EscherNet, Free3D, and One-2-3-45 on held-out objects from ObjaverseXL, as well as unseen real-world objects from Google Scanned Objects, Amazon Berkeley Objects, and the Digital Twin Catalog.
@@ 1 Introduction
Recovering 3D appearance from a single image is a hard, underspecified problem [13, 29]. A process that generates multiple views from a monocular image of a subject can serve as a useful precursor for 3D reconstruction [31, 12, 27]. Such a multiview prior is commonly implemented as a diffusion denoising probabilistic model [17] to capture a distribution of valid 3D outputs. A common pitfall, however, is that most implementations conflate the prediction of the unseen geometry and texture, making the output space complex and the task harder.
@@
Our hierarchical approach is analogous to the unCLIP approach behind DALL-E 2 [39], a popular model for text-to-image synthesis: they first map CLIP text embeddings to image embeddings using a probabilistic prior, then decode the image embeddings to pixels. Our work is not only inspired by unCLIP’s two-stage generative process, but can also be seen as downstream of unCLIP. We call our approach “unPIC” – undo-a-picture with Probabilistic Inverse Cameras – to evoke the inverse of photography.
To specify the target shape of an object before filling in appearance-level details, we use a pointmap-based intermediate representation. We adapt the idea of a Normalized Object Coordinate Space (NOCS [45]), a function that maps every point on an object’s surface to a unique RGB color based on its spatial coordinates. When rendered as images (aka pointmaps), NOCS establishes point-to-point correspondence across viewpoints (Fig. 1). It is naturally suited to help coordinate the generation of consistent novel-view images. We introduce a version of NOCS that exploits an available degree of freedom—the orientation of the color space—to make the pointmaps more predictable for the prior module. Our version, called “CROCS”, enables camera-pose-invariant predictions, which in turn enables in-the-wild predictions.
Figure 1: Top: A hierarchical approach to novel-view synthesis. A prior models multiview geometric features from a single image and are jointly decoded to the target novel-view images. Our intermediate features, CROCS, establish point-to-point correspondence across views. Bottom: Samples from the prior and decoder. Our model exhibits transferrable shape understanding having never seen a real-world pixel.
Ultimately, our goal is to produce realistic and diverse shape and appearance completions from a single 2D image. To motivate our design, we present the following experiments: we first show that a diffusion-based appearance decoder achieves better novel-view synthesis when equipped with ground-truth NOCS pointmaps (Sec. 4.1). Next, we train a diffusion-based geometric prior to predict multiview pointmaps from a given source image (Sec. 4.2). Our camera-relative version of NOCS significantly improves the diffusion prior’s performance, achieving a 4x boost in shape prediction accuracy. Finally, we stack the separately trained prior and decoder for hierarchical novel-view synthesis from a single image (Sec. 5). We perform favorably against multiple baselines CAT3D, One-2-3-45, OpenLRM, EscherNet, and Free3D on appearance and shape prediction metrics. These results demonstrate our method’s superior shape understanding, and the benefit of geometry-driven multiview synthesis.
@@ 2 Related Work
Novel view synthesis: Optimization-based NVS methods such as Neural Radiance Fields [32] or Gaussian Splatting [26] fit a geometric representation to a specific scene. Such methods need several consistent input images to be able to interpolate between camera poses. When only one image is available, we need a learned prior to predict what an object/scene looks like from alternative views. One line of work builds on spatially arranged implicit representations—such as convolutional feature grids in pixelNeRF [54] or attention-based triplane tokens in LRM [18]—to enable feed-forward inference. These original methods are deterministic, thus suffering from averaged/blurry reconstructions, as well as inconsistent 3D outputs from different views of the same object. Probabilistic extensions such as DMV3D [53] help instantiate multiple versions of an object from a single image. In a similar vein, methods such as DreamFusion [37], Score Jacobian Chaining [46], Zero-1-to-3 [31], and ZeroNVS [41] use diffusion-based priors to iterate between novel view synthesis and 3D reconstruction. Such methods tend to be slow due to scene-specific optimization (typically NeRF- or SDF-based). They also suffer from issues [2] due to bootstrapping from pretrained text-to-image models.
Recently, it has become common to generate multiple novel views before reconstructing 3D. EscherNet [27], Free3D [57], ReconFusion [50], and CAT3D [12] all take this approach, but avoid using or producing any geometric explanation during the multiview synthesis. Although generating multiple images jointly helps boost consistency on its own, the lack of geometric grounding/supervision makes these methods susceptible to producing plausible images but implausible geometry. And while diffusion-based methods like SODA’s latent bottleneck [21] and 3DiM’s stochastic conditioning [49] can help reduce inconsistencies among generated novel views, they lack explicit geometric control. Our approach allows direct geometric supervision, and ensures that generated novel views are synchronized via an explicit geometric input. At inference time, our method does not depend on having access to any geometric information. We show our method achieves better consistency among novel views relative to geometry-free approaches.
@@
Geometry-conditioned generative models: ControlNet [55] popularized the use of pixel-aligned hints to condition and guide a diffusion model. Other applications of geometry conditioning include motion brush animation [34] and pose-conditioned multi-object generation [22, 51]. This approach is especially tempting as diffusion models can be trained with conditioning dropout to enable classifier-free guidance [16], i.e., model a conditional and unconditional distribution simultaneously. At inference time, the model can cope with not having any geometric information, and still sample from the unconditional or source image-conditioned distribution. Few prior works (with the prominent exception of Motion-I2V [43]) have explored sampling geometric features using a diffusion prior, and feeding them to a decoder in place of true features or user inputs. We show this is not only possible, but desirable. Sampling from a prior ensures the geometric conditioning is realistic for the given example, diverse (e.g., different 3D shapes corresponding to a 2D image), and obtainable without human effort.
Pointmaps: Point-based representations are popular in 3D modeling. Recently, DUSt3R [47] introduced a novel approach to 3D reconstruction, estimating dense pointmaps to unproject a collection of images to 3D. DUSt3R and our work share two similarities: First, DUSt3R estimates pointmaps for two images at once to establish their correspondence. Likewise, we generate 8 pointmaps simultaneously to ensure they correspond with each other. Second, to handle scale ambiguity, DUSt3R normalizes pointmaps using their norm. This is similar to the scale normalization performed by NOCS. Two key differences between DUSt3R and our work are as follows: (i) We use a diffusion-based probabilistic model to handle the underspecified nature of image-to-3D. (ii) In DUSt3R, pointmaps are initially expressed in pairwise reference frames. These need to be aligned globally via optimization. In contrast, we express all pointmaps in a shared reference frame (i.e., CROCS color cube) from the outset. We go one step further—we set the orientation of the color cube based on the camera pose of the source image in each example. This makes the coloring predictable from arbitrary viewpoints, and learnable across examples.
Diffusion-based geometric priors: Pretrained diffusion models been shown to yield latents that are useful for various vision tasks including depth estimation and semantic correspondence [10, 9]. They have also been successfully trained to output dense annotations such as depth and surface normals [14, 25, 11]. Typically these models annotate monocular images rather than predict novel views. They lack geometric consistency when applied to different views of a scene. In fact, there is also recent work (called SpaRP [52]) which attempts to annotate images with NOCS-based pointmaps. The main difference with our work is that they only estimate NOCS for a given set of input images to infer their correspondence, rather than use NOCS to model the unseen geometry at novel views.
@@ 3.1 A Hierarchical Multiview Framework
unPIC is a hierarchical diffusion model that takes a 2D image and predicts a set of target novel-view images. Given a single view of an object, unPIC comprises: (i) A prior that infers representations of the subject’s 3D geometry. This inference step combines the task of annotating the source image as well as predicting representations for novel views. Together, these predictions form a loop in representation space. (ii) An appearance decoder that transforms all multiview geometry predictions into textured images of the subject.
Annotating the source image alone may be a deterministic step (e.g., if using features from a typical deterministic encoder). In contrast, predicting novel-view features is inherently non-deterministic, due to the partial information observed in a single image. There may be several valid loops that go through the same point that describes the source image (see Figure 2). Hence, a probabilistic prior is essential, as opposed to the bulk of prior work [54, 18, 23]. As for the appearance decoder—which takes predicted representations and maps them to images—that could be (i) either deterministic or stochastic; (ii) either separately conditioned on each single view or jointly on multiviews. Our proposed representation (CROCS) captures the geometry but leaves the texture of the subject entirely unknown. Hence we use a probabilistic model again.
Figure 2: Schrödinger’s cup: two sets of valid novel views, following different trajectories in representation space. The observed view does not reveal whether the cup has a handle or not.
Depending on the choice of representation, the target annotations may vary haphazardly as a function of the camera pose, or remain invariant (e.g., global semantic features such as the class). A suitable representation must ensure the loops are well structured, in that there is a learnable relationship between target views across examples. This increases the benefit of modeling multiple views simultaneously, as we will show empirically in Table 2.
3.2 Target-Camera Poses for Arbitrary Source Poses
Our aim is to generate novel views predicting an object’s appearance as it spins around its vertical axis. We want to do this for arbitrary source images. In keeping with prior work [12], we predict a small set of maximally informative target poses, thereby enabling downstream 3D reconstruction or camera interpolation.
Concretely, we predict novel views at K camera poses, relative to any given image. These correspond to rotations $\theta_{k}=2\pi(k-1)/K$ of the object about its vertical axis, which remains aligned with the vertical axis of the world. We include the source camera pose as one of the targets, $\theta_{1}=0$; this helps anchor the model’s predictions of novel views. One can also vary the camera elevation angle (as in [52]) along with the azimuthal angle, but to allow comparison with competing methods, we choose not to.
Note that using $K$ camera poses does not equate to a fixed multi-camera rig (as in [3]). Rather, we allow the source camera to vary freely on hemispheres of arbitrary size around the object. Given any source camera, our target cameras are at the same height as the source, the same distance to the object, but at deterministic rotations around the object’s vertical axis, and facing the object.
Since we define target camera poses relative to the source camera, we do not need to explicitly condition our model on camera poses. This precludes camera control (the ability to generate a novel view at an arbitrary pose).
@@ 3.3 CROCS: Camera-Relative Object Coordinates
We predict the geometry of an object before decoding its multiview appearance. To do so, we rely on NOCS, a scale-free representation of geometry. The object/scene is first uniformly scaled to fit a 3D unit cube, so all spatial coordinates are in $[0,1]$. These can therefore be interpreted as RGB colors and rendered as images, which maps naturally onto our multiview framework. While a single NOCS image serves as a dense image annotation, a set of NOCS images taken together forms a point cloud.
Unfortunately, NOCS on its own does not provide a predictable coloring of the object’s surface—any rotation of the RGB color space would still produce a valid coloring of a fixed object. We introduce a version of NOCS that is canonicalized to paint an object relative to the source camera pose (see Figure 3). Say the source camera moves through a rotation of $\theta$ about the object’s vertical axis. We correct for that by applying an SO2 rotation to the Red and Green channels—which span the NOCS ground plane—of all pointmaps for the given object. Note that this destroys any pose relations across objects (our training data does not, for instance, paint all cups’ handles with a consistent color).
We address two geometric subtleties in working with CROCS in Appendix B—the first one is about renormalizing pointmaps after a change of reference frame to ensure they are still bounded in $[0,1]$. The second one explains our choice not to canonicalize CROCS colors with respect to varying camera elevations $\phi$.
Our CROCS representation diverges from the typical NOCS formulation [45, 28] a second way: when dealing with a multi-object scene, we treat it as a single object. This is better suited for our task, as we care less about capturing the pose of individual objects, but more about the spatial arrangement and relative position of different objects.
Figure 3: Camera-Relative Object Coordinate Spaces. We show two data-points (Left and Right columns) obtained from one object. Top-left: The wireframe shows the RGB reference cube used to paint the object surface. The large camera denotes the source view, whereas the smaller cameras denote (3 of 7) novel views. Top-right: Say all camera locations are rotated by $\theta=120$ degrees around the vertical axis (the object stays fixed). Then we also rotate the color reference cube by the same degree. This ensures each camera faces the same side of the cube that it was facing prior to the cameras’ rotation. Bottom: Target CROCS images in clockwise order corresponding to the cameras above. In a given data-point, any part of the object is consistently colored across target images. Across data-points, although a given part of the object may change colors, its color is predictable based on its location(s) in the target image(s). Each target view has a consistent color bias that is learnable across examples.
3.4 Dataset
We train on a combination of object and scene assets from Objaverse [5] and the alignment-finetuning subset of ObjaverseXL [6] . We holdout 50k randomly sampled objects from ObjaverseXL for evaluation. For each asset, we pre-render images from a dense grid of camera poses (at various $\theta$, $\phi$, and $r$; see Section 3.2). We also pre-render one NOCS pointmap per image using a fixed NOCS RGB reference frame. At training time, we adapt the pre-rendered NOCS pointmaps to CROCS on demand based on which camera is chosen as the source, using rotations of the RGB space as mentioned in Section 3.3.
When rendering images, we add an elevated point light behind the camera. We ensure it stays fixed relative to the camera, rather than fixed in the scene. This ensures the model cannot use lighting cues; it treats each image the same in terms of lighting. (A small number of assets have preexisting lights. We do not remove these.).
3.5 Diffusion Model
We use a standard pixel-based Diffusion Denoising setup [17]. The prior and decoder models use an identical architecture, broadly following the UNet from [12]. We train the two models independently to maximize the geometric diversity of the prior—else it would be forced to output the most optimal representations for the decoder’s loss (e.g., hacks to copy bits of texture from the source image). Moreover, we train the models from scratch, without relying on pre-trained weights, to avoid ambiguity about where the models gain their representational and reconstruction abilities.
To denoise $K=8$ multiview images simultaneously, we find that tiling them to form a superimage works best. We tile clockwise along 4 columns and 2 rows—this ensures that adjacent tiles always correspond to adjacent camera poses in 3D (see  Figure 3). The source image is also provided to the model as a superimage, but with 7 of 8 tiles remaining empty. For $K=4$, the superimage grid size is 2x2.
We run preliminary experiments (Sections 4.1 and 4.2) at base image size 64x64, predicting either 1, 4, or 8 targets. We use a standard cosine diffusion schedule [33] to train these models. Our final model works on 256x256 base images, predicting 8 targets at a superimage resolution of 1024x512. To scale to this size, we use the interpolated, logSNR-offset schedule from Eq. 6 of [19]. The higher noise levels are necessary for diffusion to overcome the correlations at this resolution.
4 Preliminary Experiments
4.1 What features are most useful to decode novel views?

### v3

# How to Spin an Object: First, Get the Shape Right

@@ Abstract
Image-to-3D models increasingly rely on hierarchical generation to disentangle geometry and texture. However, the design choices underlying these two-stage models––particularly the optimal choice of intermediate geometric representations––remain largely understudied. To investigate this, we introduce unPIC (undo-a-Picture), a modular framework for empirical analysis of image-to-3D pipelines. By factorizing the generation process into a multiview-geometry prior followed by an appearance decoder, unPIC enables a rigorous comparison of intermediate geometry representations. Through this framework, we identify that a specific representation, Camera-Relative Object Coordinates (CROCS), significantly outperforms alternatives such as depth maps, pretrained visual features, and other pointmap-based representations. We demonstrate that CROCS is not only easier for the first-stage geometry prior to predict, but also serves as an effective conditioning signal for ensuring 360-degree consistency during appearance decoding. Another advantage is that CROCS enables fully feedforward, direct 3D point cloud generation without requiring a separate post-hoc reconstruction step. Our unPIC formulation utilizing CROCS achieves superior novel-view quality, geometric accuracy, and multiview consistency; it outperforms leading baselines, including InstantMesh, Direct3D, CAT3D, Free3D, and EscherNet, on datasets of real-world 3D captures like Google Scanned Objects and the Digital Twin Catalog.
@@ 1 Introduction
Recovering 3D appearance from a single image is a hard, underspecified problem [11, 25]. Models that provide 3D coverage by generating novel viewpoints of a given scene must meet two goals: one, maintain 3D consistency across the generated novel views; and two, realize diverse 3D instantiations of the given 2D input. Multiview diffusion models [21, 10, 59, 53], which sample multiple novel views together, have achieved considerable success at this task.
@@
Despite its well-motivated foundations, hierarchical 3D generation (image $\rightarrow$ 3D geometry $\rightarrow$ 3D appearance) remains understudied. Commercial approaches like Hunyuan3D [58] and CLAY [56] use specialized 3D geometry encoders to represent object shapes, with little analysis of design choices. Furthermore, there is limited insight into (a) the value of a hierarchical pipeline, and (b) the optimal choice of intermediate representations. To study these questions, we introduce a modular image-to-3D framework called unPIC (undo-a-Picture). We make two assumptions to facilitate systematic study: we use dense, image-like representations of 3D shape (e.g., depth, segmentation, or pointmaps), and off-the-shelf VAEs (from Stable Diffusion) to encode geometry and appearance.
We find a particular choice of intermediate representation outperforms others: Camera-Relative Object Coordinates. CROCS encodes the 3D coordinates of all points in a scene within a unit cube oriented toward the source camera. We find in Section 4.1 that (i) CROCS is an effective conditioning signal for the second-stage appearance decoder compared to alternatives such as depth, other pointmaps (NOCS), or pretrained visual representations (DINO or CLIP); and (ii) CROCS is also easier to predict for the first-stage geometry prior. Putting the two stages together, we show that unPIC using CROCS achieves (iii) superior novel-view accuracy and multiview consistency than strong multiview diffusion baselines including CAT3D, EscherNet, One-2-3-45, and Free3D (Section 4.2).
Conveniently, CROCS enables feedforward, direct 3D generation of a point cloud without a separate reconstruction stage. We show that (iv) direct 3D reconstruction via CROCS is more accurate than SoTA baselines Direct3D or InstantMesh (Section 4.3). Finally, we (v) ablate unPIC’s hierarchical formulation in Section 4.4, showing a significant benefit to predicting the shape of an object before filling in appearance-level details.
Figure 1: The unPIC architecture. A prior predicts multiview geometric features from a single source image. Our choice of features, CROCS, establishes point-to-point correspondence across views. The predicted CROCS images serve as a blueprint for the decoder module, which generates the fully textured target-view images, spinning the object around its vertical axis. We show 4 of 8 views here for clarity. All images are encoded as latents for diffusion. Once decoded, the output images can be directly converted to the vertices and vertex colors (respectively) of a 3D point cloud.
@@ 2 Related Work
2.1 Background
Classic novel view synthesis (NVS): Scene‑specific methods such as NeRF [28] and 3D Gaussian Splatting [20] fit explicit geometry but require multiple consistent input views. When only one image is available, these methods need a learned prior to predict what the scene looks like from alternative views. Such priors include pixelNeRF [54] and LRM [15]. They rely on spatially arranged representations to enable feed-forward NVS, but are deterministic, and suffer from averaged/blurry reconstructions.
Diffusion-based 3D generation: Early text-to-3D methods such as DreamFusion [32] and Score Jacobian Chaining [41] introduced the use of diffusion priors for iterative 3D generation. But they optimize per scene and can suffer from multi-face (Janus) artifacts. Subsequent image-to-3D models like Zero-1-to-3 [27] and ZeroNVS [35] train 3D-aware diffusion models from single images, enabling feed-forward inference, typically followed by an optimization step to produce a final 3D representation. This “generate then optimize” paradigm has become a dominant strategy for 3D reconstruction.
Multiview diffusion for NVS: Recent methods such as EscherNet [21], Free3D [59], ReconFusion [48], and CAT3D [10] denoise K views simultaneously to exchange information across images and promote consistency. Nevertheless, their lack of geometric grounding can result in generated images that are individually plausible but collectively represent an impossible 3D object. Efforts to mitigate this, such as 3DiM’s stochastic conditioning [46], still lack direct geometric control.
2.2 Motivation
@@
Geometry as an input: ControlNet [55] popularized the use of pixel-aligned hints to condition and guide a diffusion model. Other successful applications include motion-brush animation [29] and object-pose-conditioned scene/layout generation [16, 50]. These non-hierarchical approaches are especially tempting given that diffusion models can be trained with conditioning dropout to enable classifier-free guidance [14]. At inference time, the model can cope with not having any geometric information, and still sample from the unconditional distribution.
Geometry as an output: Pretrained diffusion models yield useful 3D information even when they are entirely appearance-focused [8, 7]. More geometry-focused models like He et al. [12], Ke et al. [19], Fu et al. [9] output dense annotations such as depth and surface normals, but typically annotate monocular images rather than predict novel views. They lack geometric consistency when applied to different views of a scene. We show that our pointmap-based representation (CROCS) facilitates explicit, 3D-consistent geometry prediction.
Hierarchical generation: unPIC decouples geometry and appearance by first predicting novel-view geometry via a dedicated prior, then conditioning an appearance decoder on the result. This is similar to Motion-I2V’s explicit motion generation to condition a motion-guided video model [37]. A hierarchical image-to-3D pipeline produces realistic, diverse 3D shapes without manual effort. A separation of concerns between the geometry prior and appearance decoder ensures that each stage remains specialized, promoting higher predictive diversity than end-to-end models.
2.3 Current Approaches
Geometry‑supervised image-to-3D. Methods such as One‑2‑3‑45 [26] (SDF via cost volumes) and InstantMesh [53] (triplane to mesh with depth/normal supervision) incorporate explicit geometric signals and return explicit surfaces. Closer to our work, Bolt3D [38] and CRM [45] use pointmaps as their geometric representation; however, geometry and appearance are predicted jointly (Bolt3D) or appearance-first (CRM), rather than unPIC’s geometry‑first approach. Two recent approaches, TRELLIS [51] and CLAY [56], also use a geometry-first approach, but rely on specialized 3D VAEs rather than our 2D-pretrained VAE (from Stable Diffusion 1.4). A separate line of work focuses on untextured mesh generation (e.g., CraftsMan3D [24], Direct3D [49]), emphasizing shape quality.
Pointmaps: DUSt3R [43] predicts dense pointmaps to establish 3D correspondence between images, but does not model unseen views. VGGT [42] proposes viewpoint‑invariant pointmaps with per‑scene normalization. SpaRP [52] uses a diffusion model to estimate NOCS pointmaps [40] for the given input views; in the single-image-to-3D setting, SpaRP would predict a single NOCS image to annotate the source view only, as the model does not explicitly predict novel-view geometry.
SweetDreamer [23] is closer to unPIC in its use of ‘Canonical Coordinate Maps’ to predict novel-view geometry, but still relies on the NOCS-assumption that “all objects within the same category adhere to a canonical orientation”; it further relies on optimization (score distillation) to produce the final 3D reconstruction. Our CROCS differs by being camera‑relative unlike SweetDreamer, and normalized consistently to $[0,1]$ unlike VGGT. CROCS yields predictable statistics across target viewpoints, and enables direct point‑cloud extraction during view synthesis.
@@ 3.1 Hierarchical generation
unPIC is a hierarchical generative model that takes a 2D image and predicts a set of target novel-view images. Given a single view of an object, unPIC comprises: (i) a geometry prior that infers representations of the object’s 3D geometry, and (ii) an appearance decoder that transforms all multiview geometry predictions into textured images of the subject. The two modules are trained independently. At inference time, we run the modules in a stacked fashion: the decoder takes the prior’s outputs and the source image as inputs.
This specialization and separation of concerns has two advantages: one, it helps maximize the diversity of our samples at each stage. If the prior and decoder were trained end to end, the prior would need to output the most optimal representations for the decoder, and those could be texture-specific rather than geometric. A second advantage arises from the fact that the appearance decoder is only ever trained with ground-truth geometry: it adheres closely to the geometric conditioning.
3.2 Multiview Diffusion
Both the prior and decoder modules are implemented with identical multiview diffusion architectures, building on recent work [48, 21, 59, 10] that showed the benefit of predicting all K target views simultaneously. We found that tiling the views to form a single superimage works well; we tile them clockwise to ensure each view is adjacent to its closest neighbors in camera space (see Fig 1). Our U-Net architecture follows [10], ensuring information exchange at two levels: between adjacent views using standard convolutions, and globally using self-attention across the superimage. Our implementation models K=8 views in a 2x4 grid. K=8 is sufficient to capture object geometry and appearance, and allows enough viewpoint overlap to teach the models to maintain multiview consistency, as further motivated in Appendix B.3. Since the target views are defined w.r.t. the source camera pose and remain fixed to produce a $360^{\circ}$ spin, our model does not require explicit camera poses as input. The prior and decoder modules are trained from scratch, exclusively on Objaverse and ObjaverseXL assets.
We follow the Latent Diffusion approach [34], using a VAE to encode the K views independently before tiling them to form a superimage. We fine-tune the Stable Diffusion 1.4 VAE separately on CROCS and RGB images from our training data under identical training conditions. While both VAEs start with similar PSNR scores ($\sim$37-38), the CROCS VAE’s score goes up to 51 after fine-tuning; the RGB VAE’s score improves slightly to 38. The CROCS VAE’s boost in performance comes at a lower KL representational cost than the RGB VAE’s (see Table 7 for detailed metrics). This is intriguing given that both VAEs started from the same pretrained checkpoint, which was never originally trained on CROCS images. We conclude that it is easier to reproduce CROCS pointmaps, which vary smoothly from pixel to pixel, than textured RGB images. This is yet another motivation for unPIC’s approach of predicting geometry before appearance.
Our full framework can be described as follows: Let $z^{g}_{k}$ denote the geometry latent for the $k^{th}$ view (encoded from a CROCS pointmap) and $z^{a}_{k}$ denote the appearance latent (encoded from an RGB image). We assume both kinds of latents are standardized by their empirical mean & standard deviation. They are then tiled to form the target superimages for geometry ($Z^{g}$) & appearance ($Z^{a}$):
$Z^{g}=\text{tile-clockwise}(\{z^{g}_{k}\}_{k=1}^{K})$
$Z^{a}=\text{tile-clockwise}(\{z^{a}_{k}\}_{k=1}^{K})$
Both the prior and decoder are conditioned on an input superimage, $X^{a}$, which is constructed from a masked set of K appearance latents. During training, we sample the index of a single source view $k^{*}\sim Uniform({1,…,K})$. The input latent for this view, $x_{k^{*}}$, is set to its ground-truth appearance latent, $z_{k^{*}}$, whereas the remaining K1K-1 latents are replaced by 0. While we train on only one input view, randomizing its position teaches the model to process arbitrary input configurations at inference time. The resulting input latents, $\{x^{a}_{k}\}_{k=1}^{K}$, are tiled clockwise to form the final superimage:
$X^{a}=\text{tile-clockwise}(\{x^{a}_{k}\}_{k=1}^{K})$
Finally, the unPIC inference pipeline can be written as:
$\hat{Z}^{g}\sim f_{\text{prior}}(\cdot|X^{a})\quad\text{and}\quad\hat{Z}^{a}\sim f_{\text{decoder}}(\cdot|X^{a},\hat{Z}^{g})\quad$
@@ 3.3 CROCS: Camera-Relative Object Coordinates
Geometry underpins unPIC’s hierarchical approach: we predict the 3D shape of an object before decoding its multiview appearance. To do so, we propose a dense pointmap representation called Camera-Relative Object Coordinates (CROCS). CROCS is a scale-free representation of geometry: the object/scene is first uniformly scaled to fit a 3D unit cube, so all spatial coordinates are in [0, 1]. These can therefore be interpreted as RGB colors and rendered as images, which maps naturally onto our multiview framework. While a single CROCS image serves as a dense annotation for a target novel view, a set of CROCS images taken together form the vertices of a 3D point cloud.
CROCS can be seen as a variation of Normalized Object Coordinates [40]. Instead of orienting the color space w.r.t. the object’s class-canonical pose, CROCS orients the scene geometry w.r.t. a given source camera. CROCS is also related to VGGT’s view-invariant pointmaps [42], but those are scaled by a per-scene normalization factor, and not actually bounded.
NOCS vs CROCS. While NOCS relies on a class-canonical orientation to establish correspondence between object instances, CROCS provides a pose-free, scene-level representation (see Figure 2). By orienting the coordinate system relative to the source camera rather than the object’s class, CROCS eliminates the need for object segmentation or identification. This camera-relative alignment ensures that color distributions remain predictable across target viewpoints—for example, the CROCS image for the source-camera view, corresponding to the front face of the CROCS cube, is colored white in the top-right-near corner and black in the bottom-left-far corner. This statistical predictability makes CROCS ideally suited for predicting novel views at regular intervals, as the model learns to associate specific camera aspects with biased color patterns.
We provide a precise mathematical description of CROCS in Appendix A.1. We also address two geometric subtleties in working with CROCS in Appendix A.2—the first one is about renormalizing pointmaps each time the source camera is moved to ensure they are still bounded in $[0,1]$. The second one explains why the CROCS cube remains axis-aligned, sitting squarely on the ground-plane, when the source camera is moved higher or lower (changing its elevation $\phi$).
Figure 2: Choice of pointmaps. Left: source images derived by moving the camera. NOCS: objects are normalized and colored individually based on a class-canonical pose. The same semantic part receives the same color across instances, e.g., a right arm is always blue. CROCS: objects in the scene are normalized and colored collectively. Each pixel encodes normalized $(x,y,z)$ coordinates. The source camera sets the orientation of the coordinate system (shown as the RGB cube, fixed across examples). Any target viewpoint around the cube has a predictable color pattern.
4 Experiments
We provide empirical results on our choice of representation in Sec 4.1; the quality of our novel-view synthesis in Sec 4.2; the accuracy of our 3D geometry in Sec 4.3, and finally an ablation of our hierarchical design in Sec 4.4.
4.1 Choice of Intermediate Representations
To motivate geometric representations and the use of CROCS specifically, we run two preliminary experiments: one to assess what features are predictive of novel-view images, and another experiment to assess what features are predictable from a single source image.